OpenStudy (anonymous):

Luke is twice the age of Sam . In three years he will be 3/4 the age of Luke . How old are Sam and Luke?

OpenStudy (princeharryyy):

let the age of sam be x; than lukes age = 2x given; after 3 years sam age = x+3, lukes age = 2x+3; according to question .... x+3 = 3/4(2x+3) => 4x+12 = 6x+9 => 3=2x=> x=1.5 years so current age of sam = x+3 => 1.5+3 = 4.5 yrs and lukes age = 2*1.5 + 3 = 6 years ans @Xomoneek

OpenStudy (princeharryyy):

?????